# Brain states analysis of EEG predicts multiple sclerosis and mirrors disease duration and burden

### v5

@@ 3.4 Disease predictor compared to clinical data
The brain states analysis technique separates with high probability healthy controls from patients with MS (figures 1 and 3). In order to explain this rather facile achievement, considering the data originate from single one-time clinical EEG measurements, we examined additional objective neurological measures such as EDSS values or plain disease duration which we hypothesized will correlate with our EEG analysis. These last two clinical measures arguably reflect disease related accumulation of neurological deficits.
Both plots in figure 3 imply a significant correlation of these two clinical measures with the brain states predictor space, which, we surmise, separates the two subject groups, the more the disease burden in the CNS increases. In fact, all patients to the far left in the ‘control’ brain states predictor space (marked green with letters from a to g) feature a short disease history of less than two years (mean 0.8 years) and cited motor and sensory deficits which may originate from lesions in the spinal cord rather than the upper brain parts. In contrast, the patients to the right (marked red) with longer disease histories (mean 6.0 years), self-reported also vision, balance and brain stem disturbances beside only motor and sensory issues in the limbs. This suggests more complex neurological deficit patterns that are likely related to an increase of lesion load in the cerebrum itself, the de-facto generator for the EEG signal.
Figure 3: Distinction of control subjects from patients with MS through brain states prediction. The disease predictor value corresponds to the rotated coordinates as seen in Fig. 1B. A rotation was performed using the principal components of the point distribution where the predictor value is considered as the primary principal component. (A) The upper graph illustrates this separability plotted against EDSS measures that were obtained in MS. Patients are shown as red or green dots while controls feature blue dots. Seven patients are found as outliers on the left side of the disease predictor axis and are marked with green characters (a-g) for further discussion under Results. A linear regression is shown as gray line. (B) The lower graph plots separability against disease duration in years. Two patients (d,f) lack data on disease duration and hence are not included here. Linear regression is shown as gray trend line.
@@ 4 Discussion
We successfully transformed a brain states hypothesis (Koenig et al., 1999; Janoos et al., 2011) into an analytic method for processing EEG data using a quantitative embedding and modeling approach for brain states and applied the new technique to separate the two subject groups of an earlier EEG study about MS (Jouzizadeh et al., 2021). We observed several beneficial aspects when brain states analysis is added in a context as described here.
@@ 4.1 Implications
First we demonstrate with an initial implementation of our novel algorithm, a two-phase processing pipeline with brain states detection and brain states analysis, that already regular clinical EEG data clearly carry the necessary information to discriminate two cohorts as shown here. This is to our knowledge a first successful demonstration of such a hypothesis-free space-agnostic data-driven full-brain raw-data EEG analysis approach for MS. Second we state that the inclusion of brain states in the analysis is sensible and significantly enhances the statistical robustness of the approach. Third we believe this brain states approach is easily generalizable to other disease conditions and importantly also for the intra-individual monitoring of serially acquired longitudinal EEG data.
The current method readily distinguished healthy from MS cases using clinical EEG data in spite of obstacles faced with : i) early stage of the software implementation; ii) swap in the original purpose of the approach from monitoring serial data to comparing subjects; iii) low information density in common clinical EEG recordings. We explain the analysis outcome with flow derangement between electrical mental processes in MS secondary to altered physical brain properties (Kisler et al., 2020; Paolicelli et al., 2021). This was newly confirmed in a recent study about disturbances of conductance velocity through white matter lesions (Sorrentino et al., 2022). We currently offer no method to quantify processing speed in brain states but we know already that certain feature combinations of high relevance carry slow delta EEG components associated with both eye-open and eyes-closed conditions that appear to drive the separation of controls from patients (see figure 2).
We speculate that even just small conductance alterations in MS will perceptibly skew the ’electric symphony’ among the many brain activities, ’handshakes’, synchronicity and interplay between resonating loops (Buzsáki, 2005), all the while the patient still compensates without visible clinical deficit. However, our concept about internal clockworks being disturbed between brain activities will remain hypothetical without having access to high-quality EEG data (Formica et al., 2025).
@@ 4.2 Significance
That this method succeeds on clinical-grade EEG data rests on two facts, in the processing of complete brain states and the characterizing of entire EEG recordings in form of comprehensive brain states sequences without cleansing noisy epochs. These two points highlight potential advantages of the method in that older less-advanced EEG databases could be sensibly re-analyzed with modern brain states analysis techniques, while collecting new EEG data in MS or other conditions could for budgetary reasons happen on affordable low-resolution EEG recording devices. Having said that, only high sensor density (Formica et al., 2025), high scan frequency and rich cognitive stimulative task conditions will leverage dense enough information in the EEG and hence contribute to deep brain states abstraction models. While clinical EEG data come in their own right only the highest-quality EEG data can deliver the fodder to study the internal syntax inside multi-hierarchical cognetom cascade models.
4.3 Comparison with existing methods
The here used EEG data were analyzed successfully with two methods, an earlier approach (Jouzizadeh et al., 2021) and the currently presented approach, as both separated the data into two groups, control subjects and patients with MS. The fundamentally distinct nature of the two methods, however, makes a comparative explanation for the similar results difficult. The former method relies on differentiating specific features found in space and EEG frequency bands inside heavily cleaned EEG excerpts (Jouzizadeh et al., 2021). The latter method distinguishes features in a space and time independent, non-localizational manner tracing diffusely distributed signal patterns inside brain states that are registered as comprehensive state sequences over the time course of an entire EEG recording. In essence, the here introduced latter brain states method raises the analytic precision when added to the former ‘baseline’ method which solely relies on frequency band features (table 1). It is apt to detect subtle and dispersed changes within the ’clockwork’ of EEG rhythms because these signal modulations are holistically picked up as identifiable and quantifiable biases encoded within brain states sequences secondary to disease related alterations in electrical brain rhythms.
The pace for new publications using EEG to monitor MS with modern methods is picking up. It includes approaches such as describing changes in brain dynamics (Gschwind et al., 2016) in microstates (Schependom et al., 2019). using quantitative EEG with frequency power analysis (Classen et al., 2019) for chronic neuropathic pain (Kisler et al., 2020), obstacle avoidance during gait (Santinelli et al., 2021), after optic neuritis (Pawlitzki et al., 2020), in gray matter lesions (Salim et al., 2021), tracking in vivo functional remyelination (Farley et al., 2019) or cognitive impairment (Kiiski et al., 2018; Paolicelli et al., 2021; Sarrias-Arrabal et al., 2021; Akbarian et al., 2024; Vlieger et al., 2024) or fatigue (Baldini et al., 2024; Pauletti et al., 2025) or motor disorder (Bardel et al., 2024a; Leodori et al., 2024) or auditory processing (Bissonnette et al., 2023), measuring event-related desynchronization latency (Carrere et al., 2025; Martell et al., 2025), advanced analysis methods (Mohseni and Moghaddasi, 2022; Şaşmaz Karacan and Saraoğlu, 2024; Aslan and Akşahin, 2025) with MEG data (Costers et al., 2021), resting state EEG in Alzheimer’s disease (Babiloni et al., 2021), in cognitive workload studies of neurodegenerative diseases (Figueroa-Vargas et al., 2020; Cano et al., 2024) and finally literature reviews about the use neurophysiological methods in MS (Hernandez et al., 2023; Bardel et al., 2024b; Cruciani et al., 2024; Ferreira et al., 2024; Formica et al., 2025).
The central distinction, however, of the here introduced brain states method versus those referenced above is that its modeling approach, though at an early state of development and implementation, suits at once a full range of features extracted during the data preparation phase, and operates in that sense by default much more comprehensively than techniques that focus on one specific signal feature.
@@ 4.4 Limitations
First the low spatial and temporal resolution of the analyzed EEG data, albeit acquired at clinical grade, is an important limiting factor whereas instead an increase in those acquisition parameters is indeed requisite for arriving at the maximal potential of the presented analysis technique. Second the two non-engaging experimental conditions, eyes-closed and eyes-open, further limit what instead a more engaging cognitive session could offer through exposing participants to standard audio-visual stimuli or to disease-specific experimental tasks which we expect will greatly improve sensitivity of the approach to subject- or disease-specific patterns in the EEG signal. Third the brain states epoch window kept for data-technical reasons at two seconds width proved to work surprisingly well but does no justice to the speed ephemeral electrical brain activities operate at. In other words, it is the acquisition of high-quality EEG data that will crucially counter such limitations and properly account for the short-lived transient nature of informative high-frequency signal elements.
@@ 4.5 Future objectives
We see the purpose of this publication in first line in introducing a novel hypothesis-free data-driven analysis method tested on EEG jointly with clinical data of patients with MS. However another raison-d’être is our desire to raise interest in the community towards functional brain data in longitudinal clinical studies in MS in the form of serial EEG measures (Classen et al., 2019; Farley et al., 2019; Schependom et al., 2019; Figueroa-Vargas et al., 2020; Jouzizadeh et al., 2021). Moreover we expect for the future an MS-specific EEG repository to foster novel quantitative neurocognitive methods on disease progression (Tur et al., 2023) and drug efficacy and to parallel in how the International Progressive MS Alliance (IPMSA) (Thompson et al., 2022) catalyzes research between academia and pharmaceutical industry for MRI data in MS (Eshaghi et al., 2021; Falet et al., 2022).
@@ 5 Conclusion
We introduced a novel hypothesis-free quantitative full-brain analysis approach and successfully separated EEG data paired with clinical evidence of patients suffering from MS from those of healthy subjects. Thinking forward, we envision quantitative cognitive imaging using serial EEG data to become a future main player for longitudinal clinical studies where functional neurobiological outcome measures are key, be that for monitoring MS or other illnesses with prolonged time courses such as Alzheimer, Parkinson or Huntington disease, schizophrenia, depression, migraine, HIV or systemic chemotherapy. However as the knowledge base with cognitive imaging techniques grows we believe that EEG will, as in the here presented case, also have a say in the diagnosis, subtyping and staging of primary and secondary neurological diseases.

### v6

@@ 3.3 Disease predictor compared to clinical data
The brain states analysis technique separates with surprising ease and high probability healthy controls from patients with MS (figures 1 and 3). Considering the data to originate from single clinical-grade EEG recordings, we therefore examined additional objective neurological measures such as EDSS values or plain disease duration which we hypothesized will correlate with our EEG analysis. These last two clinical measures arguably reflect disease related accumulation of neurological deficits.
Both plots in figure 3 imply an important correlation of these two clinical measures with the brain states predictor space, which, we surmise, distinguishes the better the two cohorts, the more disease burden accumulates in the CNS. In fact, all patients to the far left in the ‘control’ brain states predictor space (marked green with letters from a to g) feature a short disease duration of less than two years (mean 0.8 years) and cited motor and sensory deficits which may originate from lesions in the spinal cord or lower brain rather than upper brain parts. In contrast, patients to the right marked red, with a mean disease duration of 6.0 years, reported also vision, balance or brain stem disturbances beside motor and sensory issues in the limbs. Still, those patients in red to the right with a brief disease duration of 3 years or less, may in fact, we hypothesize, suffer from a longer, hidden disease duration, albeit clinically compensated, to amass enough supratentorial pathology that finally deranges their EEG patterns. We assume that an increase in lesion load in the cerebrum, the principal generator for the EEG signal, relates just vaguely and non-linearly to the emerging complex clinical neurological deficits, be that in cognition, EDSS, or then in form of clinical disease duration, or ‘true’ disease onset, an unknown entity for that matter.
Figure 3: Distinction of control subjects from patients with MS through brain states prediction. The disease predictor value corresponds to the rotated coordinates as seen in Fig. 1B. A rotation was performed using the principal components of the point distribution where the predictor value is considered as the primary principal component. (A) The upper graph illustrates this separability plotted against EDSS measures that were obtained in MS. Patients are shown as red or green dots while controls feature blue dots. Seven patients are found as outliers on the left side of the disease predictor axis and are marked with green characters (a-g) for further discussion under Results. A linear regression is shown as gray line. (B) The lower graph plots separability against disease duration in years. Two patients (d,f) lack data on disease duration and hence are not included here. Linear regression is shown as gray trend line. The source for the post-analysis correlation calculation is in the values for the brain state predictor space and the other two modalities such as the EDSS value or the disease duration in years. For the calculation itself the open-source python module linregress from scipy was appplied (Virtanen et al., 2020).
@@ 4 Discussion
We successfully transformed a brain states hypothesis (Janoos et al., 2011) into an analytic method for processing EEG data using a quantitative embedding and modeling approach for brain states and applied the new technique to separate the two subject cohorts of an earlier EEG study about MS (Jouzizadeh et al., 2021). We observed several beneficial aspects when brain states analysis is added in a context as described here.
@@ 4.1 Implications
First we demonstrate with an initial implementation of our novel algorithm, a two-phase processing pipeline with brain states detection and brain states analysis, that already regular clinical EEG data clearly carry the necessary information to discriminate two cohorts as shown here. This is to our knowledge a first successful demonstration of such a hypothesis-free space-agnostic data-driven full-brain raw-data EEG analysis approach for MS. Second we state that the inclusion of brain states in the analysis is sensible and enhances the statistical robustness of the approach. Third we believe this brain states approach is easily generalizable to other disease conditions and importantly also for the intra-individual monitoring of serially acquired longitudinal EEG data.
The current method readily distinguished healthy from MS cases using clinical EEG data in spite of obstacles faced with : i) early stage of the software implementation; ii) swap in the original purpose of the approach from monitoring serial data to comparing subjects; iii) low information density in common clinical EEG recordings. We explain the analysis outcome with flow derangement between electrical mental processes in MS secondary to altered physical brain properties (Kisler et al., 2020; Paolicelli et al., 2021; Salim et al., 2021; Wątorek et al., 2024). This was newly confirmed in recent studies about disturbances of conductance velocity (Sorrentino et al., 2022; Wagner et al., 2025) although we currently offer no method to quantify processing speed in brain states. However, we already know that certain feature combinations of high relevance carry slow EEG components associated with both eye-open and eyes-closed conditions, that appear to prominently drive, but still as part of the entire train of feature combinations, the separation of controls from patients (see figure 2).
We speculate that even just small conductance alterations in MS will perceptibly skew the ’electric symphony’ among the many brain activities, ’handshakes’, synchronicity and interplay between resonating loops (Buzsáki, 2005; Wątorek et al., 2024), all the while the patient still compensates without visible clinical deficit. However, our concept about internal clockworks being disturbed between brain activities will remain hypothetical without having access to high-quality EEG data (Formica et al., 2025; Wątorek et al., 2024).
@@ 4.2 Significance
That this method succeeds on clinical-grade EEG data rests on two facts, in the processing of complete brain states and the characterizing of entire EEG recordings in form of comprehensive brain states sequences without cleansing noisy epochs. These two points highlight potential advantages of the method in that older less-advanced EEG databases could be sensibly re-analyzed with modern brain states analysis techniques, while collecting new EEG data in MS or other conditions could for budgetary reasons happen on affordable low-resolution EEG recording devices. Having said that, only high sensor density (Formica et al., 2025; Wątorek et al., 2024), high scan frequency and rich cognitive stimulative task conditions will leverage dense enough information in the EEG and hence contribute to deep brain states abstraction models. While clinical EEG data come in their own right only the highest-quality EEG data can deliver the fodder to study the internal syntax inside multi-hierarchical cognetom cascade models.
@@ 4.3 Limitations
First, the low spatial and temporal resolution of the analyzed EEG data, albeit acquired at clinical grade, is an important limiting factor whereas instead an increase in those acquisition parameters is indeed requisite for arriving at the maximal potential of the presented analysis technique.
Second, the two non-engaging experimental conditions, eyes-closed and eyes-open, further limit what a more engaging cognitive session could offer. We expect that exposing participants to standard audio-visual stimuli or to disease-specific experimental tasks will greatly improve sensitivity of the approach to subject- or disease-specific patterns in the EEG signal.
Third, the brain states epoch window kept for data-technical reasons at two seconds width proved to work surprisingly well but does no justice to the speed at which ephemeral electrical brain activities operate. In other words, it is the acquisition of high-quality EEG data that will crucially counter such limitations and properly account for the short-lived transient nature of informative high-frequency signal elements.
Fourth, we are unaware of the medication given to study participants nor was this information conveyed prior (Jouzizadeh et al., 2021). Certain drug classes are known to modify physiological brain responses where patients with MS may indeed receive periodically benzodiazepines (Van Schependom et al., 2019) for example to control muscle spasms and sleep disturbances, though such symptoms tend to occur at later disease stages. This makes a benzodiazepine induced bias in the analysis outcome of the present study less likely in view of the short mean disease duration of 5.1 years. Similarly does the predominant delta character of the leading brain states feature combinations (see figure 2) point away from the beta frequency presence in benzodiazepine treated patients with MS (Van Schependom et al., 2019).
@@ 4.4 Future objectives
First we see the purpose of this publication in the introduction of a novel hypothesis-free data-driven analysis method tested on EEG and clinical data of patients with MS. However another reason is our desire to raise interest in the community towards functional brain data in longitudinal clinical studies in MS in the form of serial EEG measures (Kiiski et al., 2018).
We envision further improvements in the brain states detection (BSD) process through widening the spectrum of methods involved in analyzing the EEG signal. Analysis through multifractality and persistence effectively characterizes EEG recordings of patients with MS during eyes open and eyes closed conditions (Wątorek et al., 2024). We consider multifractality and information about long-range temporal correlations, beside FFT spectra and microstates, as procedural candidates that enrich the brain states feature space and enhance the design for time series analyses in future phases of this project.
Last, we hope a future MS-specific EEG repository will foster novel quantitative neurocognitive methods on disease progression and drug efficacy (Tur et al., 2023; Montobbio et al., 2025), similar to how the International Progressive MS Alliance (IPMSA) (Thompson et al., 2022) catalyzes research between academia and pharmaceutical industry for MRI data in MS (Eshaghi et al., 2021; Falet et al., 2022).
@@ 5 Conclusion
We introduced a novel hypothesis-free quantitative full-brain analysis approach and successfully separated EEG data paired with clinical evidence of patients suffering from MS from those of healthy subjects. Thinking forward, we envision quantitative cognitive imaging using serial EEG data to become a future main player for longitudinal clinical studies where functional neurobiological outcome measures are key, be that for monitoring MS or other neurological and psychiatric illnesses with prolonged time courses. However, as the knowledge base with cognitive imaging techniques grows, we believe that EEG will also play a role in the diagnosis, subtyping and staging of primary and secondary brain diseases.
@@ 8 Data availability statement
Appendix A Table of study participants (control subjects)
index cohort age sex
0 CT 21 m
1 CT 21 m
2 CT 19 m
3 CT 29 f
4 CT 19 f
5 CT 20 f
6 CT 18 f
7 CT 19 m
8 CT 18 m
9 CT 21 m
10 CT 19 m
11 CT 19 f
12 CT 21 f
13 CT 20 f
14 CT 25 f
15 CT 18 m
16 CT 18 m
17 CT 21 m
18 CT 21 m
19 CT 18 m
20 CT 19 m
21 CT 18 f
22 CT 28 m
23 CT 19 f
24 CT 18 f
25 CT 18 f
26 CT 20 f
27 CT 18 f
28 CT 20 f
29 CT 18 f
30 CT 18 m
31 CT 21 f
32 CT 19 f
33 CT 19 f
34 CT 19 m
35 CT 20 m
36 CT 19 m
37 CT 21 m
38 CT 18 m
39 CT 20 m
40 CT 27 m
41 CT 23 m
42 CT 25 f
43 CT 19 f
44 CT 19 m
45 CT 21 m
46 CT 31 m
47 CT 28 f
48 CT 25 f
49 CT 19 m
50 CT 23 f
51 CT 19 f
Table 2: Table is continued with patient data on the next page.
Appendix B Table of study participants (patients with MS)
[…]
duration EDSS symptoms code
0 MS 32 f 3 0 bs,vision,sensory
1 MS 28 f 11 7 sensory
2 MS 24 f 0 0 bs,balance,motor b
3 MS 37 f 2 1 sensory
4 MS 40 f 14 0 motor,sensory
5 MS 21 f 2 1 sensory g
6 MS 23 m 1 0 vision
7 MS 31 m 6 0 bs,sensory,motor
8 MS 29 m 5 1.5 bs
9 MS 27 m 4 2.5 motor
10 MS 27 m 0 f
11 MS 32 f 1 0 sensory,motor e
12 MS 26 f 0 7 motor a
13 MS 33 m 0
14 MS 31 f 0
15 MS 26 f 6 2 sensory,motor
16 MS 30 f 0 d
17 MS 20 f 1 4 motor c
18 MS 24 m 0 1 vision
19 MS 21 m 0
20 MS 31 m 0 2 sensory,motor
21 MS 37 m 15 4 vision
22 MS 33 m 5 4.5 vision,motor
23 MS 21 m 1 0 vision,balance
24 MS 53 m 10 6 sensory,motor
25 MS 35 f 8 5.5 sensory,motor
26 MS 40 f 3 2 vision,sensory,motor
27 MS 32 f 5 6 motor
28 MS 28 f 5 0 vision,motor,sensory
29 MS 32 f 3 0 vision,balance
30 MS 30 f 2 1 sensory
31 MS 27 f 13 4.5 vision,sensory
32 MS 26 f 0
33 MS 23 f 1 0 sensory
34 MS 32 f 6 0 sensory,motor
35 MS 25 f 2 3.5 vision,motor
Table 3: List of study participants with multiple sclerosis (MS). The ‘symptoms’ column reports neurological deficits for those patients where such clinical information was available. The related notes in the source data were of sparse and moderate quality. For simplicity, the neurological deficits were binned into 5 categories such as [ balance, bs, motor, sensory, vision ] (with ‘bs’ for brain stem). The ‘code’ column with alphabetic characters (a-g) stands for the seven patients that were found as outliers on the left side of the disease predictor axis marked in green letters (a-g) in figure 3.
Appendix C Mean spectra for EEG frequency bands
Figure 4: Graph showing mean spectra for EEG frequency bands. The average for each of the five common EEG frequency bands ($\delta$, $\theta$, $\alpha$, $\beta$ and $\gamma$) was computed across all subjects, separated for control and MS cohorts, and for the two experimental conditions. The eyes-open (EO) condition shows bars at full opacity, and the eyes-closed (EC) condition at half opacity.
[…]